# Semantic Ensemble Loss and Latent Refinement for High-Fidelity Neural Image Compression

### v1

@@ Abstract
Recent advancements in neural compression have surpassed traditional codecs in PSNR and MS-SSIM measurements. However, at low bit-rates, these methods can introduce visually displeasing artifacts, such as blurring, color shifting, and texture loss, thereby compromising perceptual quality of images. To address these issues, this study presents an enhanced neural compression method designed for optimal visual fidelity. We have trained our model with a sophisticated semantic ensemble loss, integrating Charbonnier loss, perceptual loss, style loss, and a non-binary adversarial loss, to enhance the perceptual quality of image reconstructions. Additionally, we have implemented a latent refinement process to generate content-aware latent codes. These codes adhere to bit-rate constraints, balance the trade-off between distortion and fidelity, and prioritize bit allocation to regions of greater importance. Our empirical findings demonstrate that this approach significantly improves the statistical fidelity of neural image compression. On CLIC2024 validation set, our approach achieves a 62% bitrate saving compared to MS-ILLM under FID metric.
Index Terms—  Neural Image Compression, Deep Generative Model, Perceptual Oriented
1 Introduction
Powered by deep learning, Learned Image Compression (LIC) has undergone remarkable progress in recent times, as evidenced by numerous studies [1, 2, 3, 4]. Some cutting-edge learning-based approaches now outperform traditional codecs like VVC [5], marking a promising future for LIC technologies. However, a notable challenge arises when these models, optimized for Mean Squared Error (MSE) metrics, produce unsatisfactory visual artifacts at lower bitrates. These artifacts, manifesting as blur and lost textures, result from a misalignment between the optimization objectives and the complexities of human visual perception. Additionally, human observers tend to focus on regions of higher salience or relevance in an image, such as faces, distinct objects, or text, while overlook less critical backgrounds or blurred sections. Thus, uniform bit allocation across an image can adversely affect the perceptual quality of salient areas, especially under limited bandwidth.
Previous methods have addressed image compression challenges by incorporating perceptual and adversarial losses [6, 7, 8, 9]. In this work, we enhance the statistical fidelity of reconstructed images by proposing a semantic ensemble loss and latent refinement. The contributions are summarized as follows:
@@ 1 Introduction
  • We refine the latent representations using semantic ensemble loss coupled with stochastic Gumbel annealing. This process can extract latent codes that comply with bit-rate constraints, balance the tradeoff between distortion and fidelity, and allocate more bits to salient image regions.
@@
2 Related Work
2.1 Learned Image Compression
Learning-based lossy image compression is predominantly formulated as a Variational Auto-Encoder (VAE) framework, as explored by Ballé et al. [10]. In this framework, an input natural image $\bm{x}$ is transformed by an analysis transform $g_{a}$ into compact latent representations $\bm{y}$. These latent variables $\bm{y}$ are then quantized into discrete values $\bm{\hat{y}}$ and compressed losslessly using entropy coding techniques. The estimated distribution of these quantized values, $p_{\hat{\bm{y}}}(\hat{\bm{y}})$, employs entropy models, such as [10, 4]. At the decoder side, the received quantized latent variables $\bm{\hat{y}}$ are transformed back into a reconstruction $\bm{\hat{x}}$ via a synthesis transform $g_{s}$. The process is formulated as:
$\bm{y}$ $=g_{a}(\bm{x};\bm{\varphi}),$ (1)
$\bm{\hat{y}}$ $=Q(\bm{y}),$
$\bm{\hat{x}}$ $=g_{s}(\bm{\hat{y}};\bm{v}),$
where $\bm{\varphi},\bm{v}$ denote parameters of encoder and decoder respectively. To bridge the discretization and amortization gaps, Yang et al. [11] introduced a semi-amortized autoencoder. This model employs a Stochastic Gumbel Annealing (SGA) technique to fine-tune the latent codes, enhancing the representations. Building on this, Gao et al. [12] proposed ’code editing’, a method that optimizes the quantization steps as a variable, enabling adaptable bitrate control and ROI-focused bit distribution.
2.2 Generative Image Compression
Adversarial loss has significantly enhanced perceptual quality in end-to-end image compression, as demonstrated in several studies [6, 7, 8, 9]. By integrating adversarial loss into the rate-distortion framework, it transforms into a rate-distortion-perception optimization problem, theoretically analyzed by Blau et al. [13]. In references [6, 7, 8, 9], a conditional GAN ensures that the decoded image remains statistically consistent with the original input. An additional discriminator introduces adversarial loss to the rate-distortion framework, which can be formulated as:
$\mathcal{L}_{\text{adv}}=\mathbb{E}_{\hat{\bm{x}}\sim P_{\hat{\bm{X}}}}\left[-\log D(\hat{\bm{x}},\hat{\bm{y}})\right].$ (2)
Furthermore, an auxiliary discriminator loss trains the discriminator to differentiate between real and reconstructed images:
$\mathcal{L}_{\text{disc}}$ $=\mathbb{E}_{\bm{x}\sim P_{\bm{X}}}\left[-\log D(\bm{x},\hat{\bm{y}})\right]$ (3)
$+\mathbb{E}_{\hat{\bm{x}}\sim P{\hat{\bm{X}}}}\left[-\log(1-D(\hat{\bm{x}},\hat{\bm{y}}))\right].$
This extension is crucial as GANs align more closely with human visual perception, offering a comprehensive optimization objective.
3 Method
Fig. 1: Framework Overview: Our process begins with training a VAE-based compression model employing our semantic ensemble loss, which is subsequently refined using SGA and a computationally simplified loss.
3.1 Architecture
@@
3.2 Training with Semantic Ensemble Loss
@@
In this equation, $\alpha$, $\beta$, $\gamma$, and $\delta$ are hyper-parameters that balance the four respective loss components. The reconstruction loss, $\mathcal{L}_{\text{rec}}$, employs the Charbonnier loss[15] to measure fidelity between the original and reconstructed images. The perceptual loss [16], $\mathcal{L}_{\text{per}}$, is derived from the $L_{2}$ loss between feature representations of the original and reconstructed images. The style loss [17, 9], $\mathcal{L}_{\text{style}}$, calculates the $L_{2}$ loss between the Gram matrices of $16\times 16$ feature patches from the original and reconstructed images, emphasizing differences in local textures. Lastly, the adversarial loss, $\mathcal{L}_{\text{adv}}$, is associated with the discriminator’s loss.
@@
3.3 Latent Refinement for Perception
@@
4 Experiments
4.1 Settings
Our model is trained using $256\times 256$ patches extracted from the test split of the OpenImages V7 dataset [19]. We employ the AdamW optimizer with parameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$, conducting the training over 2 million iterations. Our training methodology is in line with that of Muckley et al. [14], which involves a two-stage process for generative models. Initially, we perform end-to-end training with a rate-constrained loss function, optimizing both rate and distortion losses, including reconstruction, perceptual, and style loss components. This is followed by a fine-tuning stage of the decoder to incorporate an adversarial loss using a non-binary discriminator. The hyperparameters in Eq. 6 are set to $\alpha=10,\beta=1,\gamma=80,\delta=0.008$ respectively during training.
For benchmarking, we compare our results with several state-of-the-art models, including MS-ILLM [14], and leading entries from the CLIC2022 challenge, such as PO-ELIC [9], PKU-SZ [20], and IMCL-T2 [21]. We obtain reconstructed images from the official CLIC2022 challenge website and calculate metrics following the methodology described in Muckley et al. [14].
For evaluation, we use the validation set of CLIC2024 (test set of CLIC2022), comprising 30 high-resolution 2k images. Performance is assessed using reference metrics—such as DISTS [22]—and non-reference metrics, including FID [23] and KID [24]. We observed that in the CLIC2022 results, the rankings under DISTS, FID, and KID metrics align more closely with human evaluation compared to other metrics like PSNR, MS-SSIM, and LPIPS. This observation underpins our decision to focus on these three indicators for comparison. According to Muckley et al.[14] and He et al.[9], compression models optimized for traditional metrics such as MS-SSIM and MSE often yield blurring artifacts. In contrast, metrics computed in feature space, like DISTS, emphasize textural fidelity, while FID and KID, which assess statistical fidelity, are commonly used in image generation to evaluate the realism of synthesized images.
Fig. 2: Qualitative comparison of refining latent representations using different sets of hyperparameters for loss.
Fig. 3: Qualitative comparison of refining latent representations using different sets of hyperparameters for ROI-based loss.
4.2 Main Performance
@@
4.3 Qualitative Results
Fig. 2 illustrates our model’s capability to balance distortion and fidelity by adjusting the weights in Eq. 9. We present variants L0, L1, and L2, configured with $\alpha=(10,30,100)$ and $\gamma=(80,40,10)$, respectively. These adjustments demonstrate our model’s flexibility in achieving the desired balance between distortion and fidelity. Increasing $\gamma$ and reducing $\alpha$ shifts the emphasis towards fidelity, resulting in images that better match the statistical distribution of the original, as evidenced by improved FID scores. In contrast, increasing $\alpha$ and decreasing $\gamma$ focuses on reducing distortion, as reflected in higher PSNR metrics.
@@ 4.3 Qualitative Results
4.4 Ablation Study
Our ablation studies aimed to assess the impact of different components in the proposed semantic ensemble loss. We trained models with varying losses at 0.15 bpp and evaluated their performance. The results, presented in Tab. 1, indicate that omitting style loss, perceptual loss, and adversarial loss results in FIDs of 12.44, 13.32, and 13.85, respectively. These findings highlight the significance of each component in the semantic ensemble loss, particularly the non-binary adversarial loss, which proves to be the most influential.
Table 1: Ablation studies on different terms in semantic ensemble loss.
@@ 4.4 Ablation Study
5 Conclusion
This study introduces a codec optimized for visual quality, leveraging semantic ensemble loss and latent representation refinement. Our codec integrates semantic ensemble loss—comprising Charbonnier loss, perceptual loss, style loss, and non-binary adversarial loss—to enhance the model’s ability to generate finer details. To achieve bit allocation, and a balanced trade-off between fidelity and distortion, we employed SGA with learnable quantization for fine-tuning latent representations. Extensive experiments on the CLIC2024 validation dataset demonstrate that our method sets a new benchmark in image compression, excelling in perceptual quality compared to prior works.

### v2

@@ Abstract
Recent advancements in neural compression have surpassed traditional codecs in PSNR and MS-SSIM measurements. However, at low bit-rates, these methods can introduce visually displeasing artifacts, such as blurring, color shifting, and texture loss, thereby compromising perceptual quality of images. To address these issues, this study presents an enhanced neural compression method designed for optimal visual fidelity. We have trained our model with a sophisticated semantic ensemble loss, integrating Charbonnier loss, perceptual loss, style loss, and a non-binary adversarial loss, to enhance the perceptual quality of image reconstructions. Additionally, we have implemented a latent refinement process to generate content-aware latent codes. These codes adhere to bit-rate constraints, and prioritize bit allocation to regions of greater importance. Our empirical findings demonstrate that this approach significantly improves the statistical fidelity of neural image compression.
Index Terms:
Generative Image Compression, Learned Image Compression
I Introduction
Powered by deep learning, Learned Image Compression (LIC) has undergone remarkable progress in recent times, as evidenced by numerous studies [3, 2, 14, 23, 22, 8]. Some cutting-edge learning-based approaches now outperform traditional codecs like VVC [4], marking a promising future for LIC technologies. However, a notable challenge arises when these models, optimized for Mean Squared Error (MSE) metrics, produce unsatisfactory visual artifacts at lower bitrates. These artifacts, manifesting as blur and lost textures, result from a misalignment between the optimization objectives and the complexities of human visual perception. Additionally, human observers tend to focus on regions of higher salience or relevance in an image, such as faces, distinct objects, or text, while overlook less critical backgrounds or blurred sections. Thus, uniform bit allocation across an image can adversely affect the perceptual quality of salient areas, especially under limited bandwidth.
Previous methods have addressed image compression challenges by incorporating perceptual and adversarial losses [1, 17, 7, 9, 15]. In this work, we enhance the statistical fidelity of reconstructed images by proposing a semantic ensemble loss and latent refinement. The contributions are summarized as follows:
@@ I Introduction
  • We refine the latent representations using semantic ensemble loss coupled with stochastic Gumbel annealing. This process can extract latent codes that comply with bit-rate constraints, and allocate more bits to salient image regions.
@@
II Method
Figure 1: Framework Overview: Our process begins with training a VAE-based compression model employing our semantic ensemble loss, which is subsequently refined using SGA and a computationally simplified loss.
II-A Architecture
@@
II-B Training with Semantic Ensemble Loss
@@
In this equation, $\alpha$, $\beta$, $\gamma$, and $\delta$ are hyper-parameters that balance the four respective loss components. The reconstruction loss, $\mathcal{L}_{\text{rec}}$, employs the Charbonnier loss[13] to measure fidelity between the original and reconstructed images. The perceptual loss [25], $\mathcal{L}_{\text{per}}$, is derived from the $L_{2}$ loss between feature representations extracted by VGG of the original and reconstructed images. The style loss [20, 9], $\mathcal{L}_{\text{style}}$, calculates the $L_{2}$ loss between the Gram matrices of $16\times 16$ feature patches from the original and reconstructed images, emphasizing differences in local textures. Lastly, the adversarial loss, $\mathcal{L}_{\text{adv}}$, is associated with the discriminator’s loss.
@@
II-C Latent Refinement for Perception
@@
III Experiments
III-A Settings
Our model is trained using $256\times 256$ patches extracted from the test split of the OpenImages V7 dataset [12]. We employ the AdamW optimizer with parameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$, conducting the training over 2 million iterations. Our training methodology is in line with that of Muckley et al. [18], which involves a two-stage process for generative models. Initially, we perform end-to-end training with a rate-constrained loss function, optimizing both rate and distortion losses, including reconstruction, perceptual, and style loss components. This is followed by a fine-tuning stage of the decoder to incorporate an adversarial loss using a non-binary discriminator. The hyperparameters in Eq. 3 are experimentally set to $\alpha=10,\beta=1,\gamma=80,\delta=0.008$ respectively during training.
For benchmarking, we compare our results with several state-of-the-art models, including MS-ILLM [18], and leading entries from the CLIC2022 challenge, such as PO-ELIC [9], PKU-SZ [16], and IMCL-T2 [19]. We obtain reconstructed images from the official CLIC2022 challenge website and calculate metrics following the methodology described in Muckley et al. [18]. For evaluation, we use the validation set of CLIC2024 (test set of CLIC2022), comprising 30 high-resolution 2k images. Performance is assessed using reference metrics—such as DISTS [5] (better than PSNR, MS-SSIM in measuring visual consistency)—and non-reference metrics, including FID [10] and KID [21].
Figure 4: Visual comparisons using different methods at the same bitrate.
Figure 5: Qualitative comparison of refining latent representations using different sets of hyperparameters for ROI-based loss.
III-B Main Performance
@@
III-C Qualitative Results
As shown in Fig. 4, at the same bitrate, visual comparisons of the compression and reconstruction images using different methods demonstrate that our approach excels in generating details and maintaining consistency with the original image. Specifically, it outperforms POELIC and MS-ILLM in preserving finer details, such as the texture around the lion’s eye and the leaf’s structure, achieving higher fidelity to the original content.
@@ III-C Qualitative Results
III-D Ablation Study
Our ablation studies aimed to assess the impact of different components in the proposed semantic ensemble loss. We trained models with varying losses at 0.15 bpp and evaluated their performance. The results, presented in Tab. I, indicate that omitting style loss, perceptual loss, and adversarial loss results in FIDs of 12.44, 13.32, and 13.85, respectively. These findings highlight the significance of each component in the semantic ensemble loss, particularly the non-binary adversarial loss, which proves to be the most influential.
TABLE I: Ablation studies on different terms in semantic ensemble loss.
@@ III-D Ablation Study
IV Conclusion
This study introduces a codec optimized for visual quality, leveraging semantic ensemble loss and latent representation refinement. Our codec integrates semantic ensemble loss to enhance the model’s ability to generate finer details. To achieve better performance and bit allocation, we employed SGA with learnable quantization for fine-tuning latent representations. Extensive experiments demonstrate that our method sets a new benchmark in image compression, excelling in perceptual quality compared to prior works.
Acknowledgment
This work was supported in part by National Key Research and Development Program of China under Grant 2022YFF1202104, in part by National Natural Science Foundation of China under Grants 62301188, 92270116 and U23B2009, in part by China Postdoctoral Science Foundation under Grant 2022M710958, and in part by Heilongjiang Postdoctoral Science Foundation under Grant LBH-Z22156.